### v2

# Exploring the Limits of Vision-Language-Action Manipulations in Cross-task Generalization

Abstract
[…]
Figure 1: The proposed AGNOSTOS benchmark evaluates cross-task zero-shot generalization through two difficulty levels. Level-1 testing involves 13 unseen tasks sharing partial similarity (objects or motions) with seen tasks. Level-2 testing has 10 unseen tasks from entirely novel scenarios, requiring stronger generalization capabilities. We systematically assess three broad categories of vision-language-action models, revealing critical limits in their ability to adapt to unseen tasks.
1 Introduction
Vision-Language-Action (VLA) models [3, 4, 5, 6, 2, 7, 8, 1, 9] have motivated a new era of robotic manipulation [10, 11, 12, 13, 14, 15, 16] by integrating visual perception, language understanding, and action generation. Through large-scale pre-training on diverse data, including human videos [17, 18, 19], real or simulated cross-embodiment robotic demonstrations [3, 20, 21], VLA models can effectively generalize across visual variations within known tasks (i.e., within-task generalization of seen tasks), such as handling objects in novel scenes or with altered properties. However, the true promise of VLAs lies in their capacity to generalize across tasks: to handle previously unseen combinations of objects, goals, and actions without prior exposure. This capability-cross-task zero-shot generalization-is essential for real-world deployment, where robots are expected to tackle novel tasks as they arise dynamically.
Despite the rapid progress in VLA research, most prior work [3, 4, 7, 8, 1] has focused on generalization testing in real-world environments. However, these evaluations are typically non-reproducible, and rarely target cross-task (i.e., unseen task) zero-shot generalization. Recently, many works [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] have been devoted to developing comprehensive simulated benchmarks, which could be utilized to evaluate the generalization of VLA models. While promising, these efforts mainly assess within-task generalization, leaving cross-task zero-shot generalization largely unexplored.
To address this critical gap, we present AGNOSTOS, a novel benchmark for evaluating cross-task zero-shot generalization in robotic manipulation. Built on RLBench [25], our benchmark comprises 23 unseen tasks that are carefully curated to differ from 18 commonly used seen training tasks [34, 35]. As shown in Figure 1, to probe different aspects of generalization, these unseen tasks are categorized into two difficulty levels:
  • Level-1 (13 tasks) shares partial semantics (e.g., similar objects like “cups” or motions like “put”) with seen tasks.
  • Level-2 (10 tasks) introduces entirely novel scenarios with no overlapping objects or actions.
We benchmark three broad categories of VLA models:
  1. 1.
    Models [1, 2, 7, 8] trained on large-scale real-world robotic demonstrations [36] or built on multimodal large language models (MLLMs);
  2. 2.
    Models [37, 38, 39] pre-trained on large-scale human action video datasets [17];
  3. 3.
    Models [40, 34, 41] trained purely on in-domain RLBench data with advanced architectures.
Our empirical findings reveal a key limitation: none of the existing models generalize effectively to unseen tasks, highlighting the need for approaches that directly address cross-task generalization.
Motivated by the success of in-context learning [42, 43, 44] in large language models (LLMs), we propose a Cross-task In-context Manipulation (X-ICM) method to address the challenges of cross-task zero-shot generalization. X-ICM uses demonstrations from seen tasks as in-context examples to prompt LLMs to generate action plans for unseen tasks. A key challenge under this setup is selecting relevant demonstrations; irrelevant prompts fail to activate appropriate knowledge in the LLM, leading to poor cross-task predictions. To address this, we introduce a dynamics-guided sample selection strategy. This approach leverages learned representations of task dynamics—captured by predicting final observations from initial states and task descriptions—to identify relevant demonstrations across tasks. Guided by the learned dynamics, the dynamics-aware prompts can be formed that enable X-ICM to elicit high-quality action predictions from LLMs, even for unfamiliar, unseen tasks.
Our contributions are threefold:
  • We introduce AGNOSTOS, the first benchmark to systematically evaluate cross-task zero-shot generalization in robotic manipulation for VLAs, with 23 unseen tasks spanning two levels of generalization difficulty.
  • We propose X-ICM, a novel method that combines in-context prompting and dynamics-guided sample selection to enhance cross-task zero-shot transfer in VLA models.
  • We conduct extensive evaluations of diverse VLA models on AGNOSTOS, uncovering fundamental limitations in current approaches and demonstrating the superior generalization of X-ICM.
2 Related Works
Vision-Language-Action Models. Vision-Language-Action (VLA) models [3, 4, 5, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56], designed to enable robots to understand instructions and interact with the physical world, have largely followed two main paradigms. The first involves modular-based approaches [2, 57, 58, 59, 60, 61, 62, 63], where different MLLM components handle perception, language understanding, planning, and action execution. For example, VoxPoser [2] uses MLLMs to synthesize composable 3D value maps for manipulation. The second paradigm focuses on end-to-end approaches [3, 36, 8], training a policy model to directly map raw sensory inputs (e.g., vision, language instructions) to robot actions. The success of these models heavily depends on the scale and diversity of training data, which comes from various data sources, including large-scale human action videos [17] and large-scale cross-embodiment robotic data [36, 64]. Based on the data, OpenVLA [7] is the first fully open-sourced work that significantly promotes the development of the VLA community. $\pi_{0}$  [1] proposes a VLM-based flow matching architecture, which is trained on a diverse dataset from multiple dexterous robot platforms.
Benchmarks for Vison-Language-Action Models. Evaluating the generalization capabilities of VLA models requires robust and comprehensive benchmarks. While existing VLAs mainly focus on real-world testing, which suffers from reproducibility issues and rarely focuses systematically on zero-shot generalization to unseen tasks. Recently, many simulated benchmarks [22, 23, 24, 25, 26, 27, 28, 30, 31, 32, 33] have been proposed to offer a controlled environment for rigorous evaluation. Colosseum [31] enables evaluation of models across 14 axes of perturbations, including changes in color and texture, etc. GemBench [32] designs four levels of generalization, spanning novel placements, rigid and articulated objects, and complex long-horizon tasks. We find that these benchmarks have predominantly concentrated on evaluating visual variations within tasks, assessing how well models can generalize to new scenes or altered attributes of objects within the same task. However, these benchmarks do not focus on the more challenging aspect of cross-task zero-shot evaluation, where models must generalize to entirely new tasks with unseen combinations of object categories and motions. This gap in evaluation limits our understanding of the true generalization capabilities of VLA models. Motivated by this, this work proposes the first cross-task zero-shot manipulation benchmark, expanding the evaluation scope of the generalization capabilities of VLAs.
In-context Learning with Large Language Models (LLMs). LLMs [65, 42, 66, 67, 68, 69, 70, 71] have demonstrated a remarkable capability known as in-context learning (ICL) [42, 43, 44], where they can learn to perform new tasks based on a few examples provided within the prompt, without requiring updates to parameters. The potential of ICL is being explored in robotics [72, 73, 74, 75, 76, 77]. Prior works like KAT [72] and RoboPrompt [76] have shown that off-the-shelf LLMs can predict robot actions directly using within-task in-context samples. By extending their paradigms to a cross-task setting, this work specifically focuses on the cross-task zero-shot generalization problem. Our X-ICM model uses demonstrations from seen tasks as in-context examples to prompt LLMs for action generation in completely unseen tasks. Under the cross-task zero-shot setting, the selection of in-context samples [78, 79, 80, 81] is crucial for a robust generalization. Our X-ICM method addresses this challenge by introducing a dynamics-guided sample selection strategy, ensuring that the selected seen demonstrations are relevant to the unseen tasks, thereby stimulating the cross-task generalization capabilities of LLMs.
3 AGNOSTOS: Cross-task Zero-shot Generalization Benchmark
To systematically assess the cross-task zero-shot generalization ability of vision-language-action (VLA) models, we introduce AGNOSTOS, a reproducible benchmark built upon the RLBench simulation environment. AGNOSTOS features 18 seen tasks for training and 23 unseen tasks for rigorous generalization testing.
Training. For training, we adopt the standard set of 18 RLBench tasks that are widely used in prior work [34, 35]. Examples of these seen tasks are shown in Figure A1. We collect 200 language-conditioned demonstrations per task, resulting in 3600 demonstrations in total. These demonstrations enable VLA models to be fine-tuned to reduce the domain and embodiment gaps between pre-training data and RLBench data.
Testing. As illustrated in Figure 1, AGNOSTOS comprises 23 held-out unseen tasks with semantics that are disjoint from the seen set (videos of all tasks are available in the Supplementary Materials). We categorize the unseen tasks into two difficulty levels. Level-1: 13 tasks that share partial semantic similarity with seen tasks—either in object types (e.g., cups) or motion primitives (e.g., stacking). Level-2: 10 tasks that exhibit no overlap in either object categories or motion types, requiring broader compositional reasoning and semantic extrapolation. Details on task curation and difficulty categorization are provided in Section A1.1 of the Appendix. For each unseen task, we perform three test runs with different seeds, each consisting of 25 rollouts, and report the mean and standard deviation of success rates.
To explore generalization boundaries, AGNOSTOS evaluates three broad families of VLA [^1] We only consider evaluating the VLA models that are fully open-source.models:
  1. 1.
    Foundation VLA models: trained on large-scale real-world cross-embodiment robotic data [36] or built upon LLM or VLM models, including OpenVLA [7], RDT [8], $\pi_{0}$  [1], LLARVA [82], SAM2Act [41], 3D-LOTUS++ [32], and VoxPoser [2].
  2. 2.
    Human-video VLA models: pre-trained on large-scale human action videos [17, 83] to capture rich human-object interactions for downstream robotic fine-tuning, including R3M [37], D4R [38], R3M-Align [39], and D4R-Align [39].
  3. 3.
    In-domain VLA models: trained from scratch on RLBench’s 18 seen tasks with task-specific model architectures. These serve as strong baselines without domain mismatch, including PerAct [34], RVT [35], RVT2 [40], Sigma-Agent [84], and Instant Policy [74].
To ensure a fair comparison, we fine-tune all models on the same 18 seen tasks when the models involve the embodiment gaps, following the official protocols of each method. A detailed description of the fine-tuning process and evaluation on AGNOSTOS is provided in Section A1.2 of the Appendix.
Table 1: Comparison of manipulation benchmarks that focuses on cross-task generalization testing. The comparison evaluates whether each benchmark includes cross-task testing, the types of VLA models evaluated, and the number of Level-1 and Level-2 unseen tasks included.
Benchmark Simulator No. of train tasks No. of test tasks Cross-task Zero-shot Generalization Test
Cross task Evaluated VLA models Level-1 unseen tasks Level-2 unseen tasks
In-domain Human-video Foundation
RLBench-18Task [34] RLBench 18 18 - - - - -
CALVIN [24] PyBullet [85] 34 34 - - - - -
Colosseum [31] RLBench 20 20 - - - - -
VLMBench [30] RLBench 8 8 - - - - -
Ravens [86] PyBullet [85] 10 10 3 0
VIMA-Bench [87] Ravens 13 17 4 0
GemBench [32] RLBench 16 44 15 0
AGNOSTOS (Ours) RLBench 18 23 13 10
Table 1 compares AGNOSTOS with existing robotic manipulation benchmarks in terms of their support for cross-task evaluation, i.e., considering test tasks that have different object categories or motions from training tasks. While many benchmarks focus on within-task visual generalization, few explicitly support cross-task zero-shot generalization test, and even fewer include tasks as challenging as our Level-2 testing scenarios. Even benchmarks that include some form of cross-task evaluation [^2] We carefully categorize their test tasks into our defined two-level difficulty task sets. often test on a narrow set of tasks, and typically evaluate only in-domain models, neglecting foundation and human-video pre-trained VLAs. AGNOSTOS fills this gap by supporting broad model coverage and introducing novel, difficult tasks in Level-2, providing a more comprehensive and diagnostic assessment of cross-task generalization.
4 X-ICM: Cross-task In-context Manipulation Method
To push the boundaries of cross-task zero-shot generalization in vision-language-action (VLA) models, we propose a method called Cross-task In-context Manipulation (X-ICM). Leveraging the cross-task generalization capabilities of LLMs, X-ICM utilizes demonstrations from seen tasks as in-context examples. The dynamic characteristics of these examples are used to prompt the LLM to predict action sequences for unseen tasks. In contrast to prior works [76, 72] that apply LLMs’ in-context learning to within-task generalization, our work is the first to extend this paradigm to a cross-task zero-shot setting. A central challenge in this setting is that the selection of in-context demonstrations significantly affects generalization performance. To address this, we design a dynamics-guided sample selection module that measures similarities between dynamic representations of seen and unseen tasks to guide the selection process, resulting in improved cross-task generalization.
Figure 2: X-ICM Method Overview. X-ICM employs a dynamics-guided sample selection module to retrieve effective demonstrations from seen tasks for each tested unseen task. These demonstrations are then used by the cross-task in-context prediction module to construct the prompt that drives the LLM to predict the corresponding action sequence.
4.1 Problem Definition and Method Overview
We tackle the problem of cross-task zero-shot robotic manipulation by exploiting the in-context learning ability of off-the-shelf LLMs. We assume access to a dataset of demonstrations from seen tasks, denoted as $\mathcal{D}^{s}=\{V_{i}^{s},A_{i}^{s},L_{i}^{s}\}_{i=1}^{N}$, where $N$ is the total number of seen demonstrations, $V^{s}_{i}=\{v^{s}_{i,1},\cdots,v^{s}_{i,T}\}$, $A^{s}_{i}=\{a^{s}_{i,1},\cdots,a^{s}_{i,T}\}$, and $L^{s}_{i}$ are the visual observation sequence, action sequence, and language description of the $i$-th seen demonstration ($T$ is the sequence length, which varies for each demonstration). For an unseen task with the given initial visual observation $v^{u}$, and language description $L^{u}$, our cross-task zero-shot manipulation method can be formulated as follows:
$A^{u}_{pred}=\{a^{u}_{1},\cdots,a_{t}^{u}\}=LLM(\mathcal{P}(\mathcal{F}^{sel}(\mathcal{D}^{s}),v^{u},L^{u})),$ (1)
where $\mathcal{F}^{sel}$ is the cross-task in-context demonstration selection process, $\mathcal{P}$ is the cross-task in-context prompt construction process, and $A^{u}_{pred}$ is the predicted action sequence for the tested unseen task.
Figure 2 outlines our X-ICM framework, which comprises two core modules. The Dynamics-guided Sample Selection module introduces a dynamics diffusion model, and then retrieves effective demonstrations based on dynamic similarities between seen and unseen demonstrations. The Cross-task In-context Prediction module constructs the LLM prompt using retrieved demonstrations to predict action sequences for the unseen task. Next, we describe each module in detail below.
4.2 Dynamics-guided Sample Selection module
Selecting effective in-context examples is essential for achieving robust cross-task generalization. We propose a two-stage Dynamics-guided Sample Selection module that learns dynamic representations from demonstrations using diffusion models, and retrieves examples that are most relevant to the current unseen task based on the similarities between the dynamic representations of demonstrations.
Diffusion-based dynamics modeling stage. To effectively capture the dynamic representations within each demonstration, we train a dynamics diffusion model $\mathcal{G}$ on all seen demonstrations. For each demonstration (e.g., the $i$-th), the dynamics diffusion model $\mathcal{G}$ takes the initial visual observation $v^{s}_{i,1}$ and language description $L^{s}_{i}$ as inputs, and predicts the future observation that matches the final visual observation $v^{s}_{i,T}$. The model $\mathcal{G}$ is initialized from InstructPix2Pix [88] and is optimized by:
$\min_{\mathcal{G}}\mathbb{E}_{i,z,\epsilon\sim\mathcal{N}(0,I)}\left[\left\|\epsilon-\epsilon_{\mathcal{G}}(v^{s}_{i,T,z},z,v^{s}_{i,1},L^{s}_{i})\right\|_{2}^{2}\right],$ (2)
where $z$ is the diffusion timestep, $\epsilon$ is the noise, $v^{s}_{i,T,z}$ is the noised final observation, and $\epsilon_{\mathcal{G}}$ is the noise predictor in model $\mathcal{G}$. By predicting the future, the inherent dynamics of each demonstration can be effectively modeled. Figure A5 shows the generation results of our dynamics diffusion model.
Dynamics-guided retrieving stage. Once trained, we use $\mathcal{G}$ to extract a dynamic feature $f^{s}_{i}$ for the $i$-th seen demonstration:
$f^{s}_{i}=[f^{s,vis}_{i},f^{s,lang}_{i}]=\mathcal{G}(v^{s}_{i,1},L^{s}_{i})\in\mathbb{R}^{2\times 1024},$ (3)
where $f^{s,vis}_{i}\in\mathbb{R}^{1024}$ is the predicted latent feature of the target visual observation, $f^{s,lang}_{i}\in\mathbb{R}^{1024}$ is the textual feature of the language description. For the unseen task, we similarly compute its feature $f^{u}$. Then, we compute cosine similarities between features and select the top-$K$ seen demonstrations with the highest similarity:
$R_{i}$ $=\frac{f^{u}\cdot f^{s}_{i}}{\|f^{u}\|\|f^{s}_{i}\|}\quad i\in\{1,\dots,N\},$ (4)
$\mathcal{I}_{idx}$ $=\operatorname{arg\,topk}_{i\in\{1,\dots,N\}}^{K}(R_{i}),$ (5)
where $R_{i}$ represents the cross-task dynamics similarity score between the tested unseen task and the $i$-th seen demonstration. And the set $\mathcal{I}_{idx}$ contains the indices of $K$ seen demonstrations that achieve the highest similarity scores. These $K$ seen demonstrations are retrieved to construct the cross-task in-context prompt.
4.3 Cross-task In-context Prediction module
For a tested unseen task, we use the obtained $K$ most-relevant seen demonstrations to construct the in-context prompt for cross-task action prediction with LLMs. Our Cross-task In-context Prediction module follows the practices in existing within-task in-context manipulation methods [76, 72], but we are the first to extend this paradigm to the cross-task zero-shot setting.
Textualizing key information. As shown at the bottom of Figure 2, for each selected seen demonstration, we extract the 3D positions of the centers of objects and the key-action sequence. For the 3D positions of objects’ centers, we use GroundingDINO [89] to detect the 2D positions of objects, and then use the depth, intrinsic, and extrinsic information of the camera to obtain the 3D coordinates of objects’ centers under the robot-base coordinate system. Additionally, we evenly divide the workspace into $100\times 100\times 100$ grids, and then normalize the 3D coordinates of the objects’ centers into the grid’s coordinate system. Thus, for the $m$-th object $O_{k}^{m}$ in the $k$-th seen demonstration, we textualize it as $O_{k}^{m}=$$objectname$: [$x$, $y$, $z$]” (e.g., “block: [52, 50, 19]”). For the key-action sequence, following [90], we select the actions when the gripper state changes or the joint velocities are near zero. In this way, the execution process of a demonstration can be determined by several key-actions. We use the 7-dimensional state of the end-effector to represent the robot action, i.e., the 3D positions, the roll-pitch-yaw angles, and the gripper’s open/close. The 3D positions of the end-effector are also normalized into the grid’s coordinate system of the workspace. And the angles of end-effector are discretized into 72 bins, with each occupying 5 degrees. Thus, for the $t$-th key-action $A_{k}^{t}$ in the $k$-th seen demonstration, we textualize it as $A_{k}^{t}=$ “[$x$, $y$, $z$, $roll$, $pitch$, $yaw$, $gripper$]” (e.g., “[53, 57, 17, 0, 36, 53, 0]”).
Constructing in-context prompt. With the above textualization of object and key-action information, for each seen demonstration (e.g., the $k$-th), we can textualize it as the mapping from its language description and object context to its key-action context:
[…]
We textualize all $K$ selected seen demonstrations in this manner, and concatenate them in descending order based on their dynamics similarity scores to the tested unseen task. For the tested unseen task, we textualize it using its language description and object context, i.e., “$[L^{unseen},\{O^{1},\cdots,O^{m},\cdots\}]\to$”. Finally, the cross-task in-context prompt is formed by concatenating the system prompt, the textualized all seen demonstrations, and the textualized tested unseen task (see the bottom of Figure 2 and an example in Figure A6 of the Appendix). This prompt will serve as the input of LLMs, which incentivizes the cross-task generalization capability of LLMs to predict the key-action sequence for the tested unseen task.
5 Experiments
Implementation Details. For our X-ICM method on the AGNOSTOS benchmark, we use a total of $N=3600$ seen demonstrations. During in-context prompt construction, we select $K=18$ demonstrations. We mainly use off-the-shelf Qwen2.5-Instruct [70] models with 7B and 72B parameters, referred to as X-ICM (7B) and X-ICM (72B), respectively. These are deployed using two or eight A6000 GPUs. For a fair comparison with existing zero-shot baselines (e.g., VoxPoser [2]), in simulation we use the ground-truth positions of objects. Ablation on using different sizes of LLMs is presented in Sec A2.2 of the Appendix.
5.1 Benchmarking Vision-Language-Action Models
Table 2 presents the cross-task zero-shot performance of our X-ICM models on 23 unseen tasks from the benchmark, including 13 Level-1 and 10 Level-2 tasks. These tasks are designed to evaluate the generalization ability of VLA models under varying levels of difficulty. We compare our models against a diverse set of baselines, including models trained on in-domain RLBench data, human videos, and LLM/VLM-based foundations. Key observations include:
  • X-ICM (7B) and X-ICM (72B) achieve average success rates of 23.5% and 30.1%, respectively, outperforming all existing VLA models.
  • On Level-1 tasks, X-ICM (7B) surpasses the prior SoTA $\pi_{0}$  [1] by 6.9%. For Level-2 tasks, performance gains are more pronounced with the 72B model.
  • While some VLA foundation models show decent performance, particularly $\pi_{0}$ on Level-1 tasks and SAM2Act [41]/3D-LOTUS++ [32] on Level-2 tasks, they fall short of the broad generalization capabilities demonstrated by our X-ICM models.
  • All prior models completely fail on at least eight of the 23 tasks. In contrast, X-ICM (7B) fails on only two, and X-ICM (72B) succeeds on all.
Table 2: Cross-task zero-shot manipulation performance on 23 unseen tasks, where the column headers show the abbreviation of each unseen task (see full task names in Table A1). N/A indicates that the tested tasks are removed since they overlap with the training tasks of the methods. Level-1 and Level-2 represent tasks with two different difficulty levels. The prefix * indicates second best.
methods Level-1 tasks
Toilet Knife Fridge Microwave Laptop Phone Seat LampOff LampOn Book Umbrella Grill Bin
in-domain training PerAct [34] 0.0 5.3 37.3 64.0 2.7 0.0 72.0 0.0 1.3 0.0 1.3 8.0 54.7
RVT [35] 0.0 2.7 50.7 26.7 50.7 2.7 40.0 0.0 1.3 0.0 1.3 0.0 6.7
Sigma-Agent [84] 0.0 9.3 56.0 9.3 30.7 1.3 65.3 1.3 0.0 0.0 0.0 1.3 4.0
RVT2 [40] 0.0 1.3 0.0 17.3 42.7 1.3 62.7 2.7 1.3 0.0 1.3 5.3 34.7
InstantPolicy [74] 0.0 1.3 13.3 4.0 4.0 18.7 24.0 0.0 0.0 0.0 0.0 0.0 0.0
human-video pretraining D4R [38] 0.0 8.0 32.0 30.7 24.0 0.0 65.3 20.0 4.0 0.0 0.0 0.0 0.0
R3M [37] 0.0 0.0 37.3 22.7 25.3 1.3 62.7 6.7 4.0 0.0 0.0 0.0 0.0
D4R-Align [39] 0.0 2.7 45.3 74.7 24.0 0.0 41.3 0.0 0.0 1.3 0.0 0.0 0.0
R3M-Align [39] 0.0 4.0 49.3 25.3 21.3 0.0 49.3 0.0 5.3 0.0 0.0 1.3 1.3
VLA foundations OpenVLA [7] 0.0 5.3 38.7 40.0 57.3 0.0 53.3 12.0 1.3 1.3 0.0 10.7 0.0
RDT [8] 0.0 0.0 46.7 13.3 14.7 0.0 50.7 0.0 0.0 1.3 0.0 8.0 0.0
$\pi_{0}$
 [1]
0.0 5.3 85.3 24.0 40.0 1.3 64.0 18.7 8.0 1.3 0.0 33.3 1.3
LLARVA [82] 0.0 0.0 12.0 0.0 6.7 0.0 40.0 0.0 0.0 0.0 0.0 0.0 0.0
3D-LOTUS [32] 0.0 6.7 N/A N/A N/A 0.0 6.7 0.0 0.0 0.0 0.0 13.3 5.3
3D-LOTUS++ [32] 0.0 5.3 N/A N/A N/A 9.3 68.0 10.7 0.0 0.0 0.0 29.3 13.3
SAM2Act [41] 0.0 0.0 36.0 40.0 6.7 6.7 62.7 6.7 0.0 1.3 1.3 9.3 0.0
VoxPoser [2] 0.0 0.0 0.0 0.0 5.3 8.0 28.0 88.7 25.3 0.0 0.0 0.0 82.7
Ours X-ICM (7B) 1.3 26.7 22.7 45.3 33.3 57.3 48.0 58.7 50.7 1.3 0.0 8.0 18.7
X-ICM (72B) 6.7 69.3 12.7 58.7 34.0 68.0 51.3 86.7 74.7 2.0 1.3 5.3 18.7
methods Level-2 tasks Level-1 Level-2 All
USB Lid Plate Ball Scoop Rope Oven Buzz Plants Charger avg (std) avg (std) avg (std)
in-domain training PerAct [34] 58.7 2.7 0.0 0.0 0.0 0.0 1.3 4.0 6.7 2.7 19.0 (1.4) 7.6 (1.1) 14.0 (0.9)
RVT [35] 89.3 2.7 0.0 0.0 0.0 0.0 4.0 8.0 5.3 4.0 14.0 (1.4) 11.3 (1.6) 12.8 (0.2)
Sigma-Agent [84] 88.0 0.0 0.0 0.0 0.0 0.0 4.0 8.0 5.3 1.3 13.7 (1.6) 10.7 (1.7) 12.4 (0.4)
RVT2 [40] 22.7 40.0 0.0 0.0 0.0 0.0 0.0 1.3 1.3 1.3 13.1 (0.4) 6.7 (1.3) 10.3 (0.6)
InstantPolicy [74] 26.7 1.3 0.0 0.0 0.0 0.0 0.0 1.3 0.0 0.0 4.3 (4.2) 2.9 (1.4) 3.7 (3.0)
human-video pretraining D4R [38] 98.7 0.0 0.0 0.0 0.0 0.0 1.3 1.3 1.3 4.0 14.1 (0.3) 10.7 (0.2) 12.6 (0.2)
R3M [37] 48.0 0.0 0.0 0.0 0.0 0.0 8.0 2.7 2.7 1.3 12.3 (1.4) 6.3 (0.9) 9.7 (0.6)
D4R-Align [39] 89.3 1.3 0.0 0.0 0.0 0.0 8.0 6.7 0.0 1.3 14.5 (1.0) 10.7 (0.2) 12.8 (0.6)
R3M-Align [39] 90.7 0.0 1.3 0.0 0.0 0.0 2.7 13.3 4.0 0.0 12.9 (0.7) 11.2 (0.7) 12.2 (0.3)
VLA foundations OpenVLA [7] 77.3 0.0 0.0 0.0 0.0 0.0 6.7 5.3 2.7 0.0 16.9 (1.3) 9.2 (0.7) 13.6 (0.8)
RDT [8] 100.0 29.3 4.0 0.0 0.0 0.0 8.0 2.7 0.0 0.0 10.4 (0.5) 14.4 (0.9) 12.1 (0.4)
$\pi_{0}$
 [1]
97.3 0.0 1.3 0.0 0.0 0.0 14.7 5.3 1.3 0.0 *21.7 (0.4) 12.0 (0.9) *17.5 (0.4)
LLARVA [82] 24.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 4.5 (0.1) 2.4 (0.0) 3.6 (0.1)
3D-LOTUS [32] 85.3 0.0 1.3 0.0 0.0 0.0 0.0 4.0 0.0 0.0 3.2 (0.5) 9.1 (0.7) 6.2 (0.5)
3D-LOTUS++ [32] 90.7 30.7 0.0 0.0 5.3 1.3 8.0 8.7 6.7 0.0 13.6 (1.0) 15.1 (1.1) 14.4 (1.0)
SAM2Act [41] 92.0 49.3 0.0 0.0 0.0 0.0 1.3 6.7 4.0 5.3 13.1 (0.4) *15.9 (1.3) 14.0 (0.7)
VoxPoser [2] 32.0 76.0 0.0 8.0 0.0 0.0 0.0 1.3 0.0 4.0 18.1 (0.4) 12.1 (0.4) 15.6 (0.2)
Ours X-ICM (7B) 98.7 20.0 6.7 9.3 0.0 6.7 16.0 2.7 5.3 4.0 28.6 (1.9) 16.9 (1.3) 23.5 (1.6)
X-ICM (72B) 98.7 13.3 4.7 36.0 0.7 16.0 20.7 7.3 2.7 2.7 37.6 (1.4) 20.3 (1.7) 30.1 (1.0)
Table 3: Effects of dynamics-guided sample selection module and different model sizes.
Models Level-1 Level-2 All
X-ICM (72B) w/o sel 30.7 (4.7) 18.0 (2.2) 25.2 (3.2)
X-ICM (72B) 37.6 (1.4) 20.3 (1.7) 30.1 (1.0)
Table 4: Effect of using different LLM models.
LLMs Level-1 Level-2 All
Deepseek-R1-Distill-Qwen-7B 10.7 (1.1) 7.9 (0.5) 9.5 (0.5)
Llama3.0-8B-Instruct 17.4 (0.7) 11.7 (1.6) 15.1 (0.3)
Ministral-8B-Instruct-2410 22.9 (0.7) 14.8 (0.3) 19.5 (0.5)
InternLM3-8B-Instruct 27.9 (0.7) 13.3 (0.4) 21.8 (0.5)
Qwen2.5-7B-Instruct 28.6 (1.9) 16.9 (1.3) 23.5 (1.6)
Figure 3: The effect of varying the number of in-context demonstrations.
5.2 Ablation Studies
Dynamics-guided sample selection. We assess the impact of the dynamics-guided sample selection module by comparing X-ICM (72B) with and without it (denoted as w/o sel, using random sampling) at the top-left of Table 3. Incorporating the module notably boosts performance and robustness across both task levels. This confirms its effectiveness in identifying informative demonstrations that enhance the LLM’s cross-task generalization ability. Further analysis, including visualizations from the dynamics diffusion model and comparisons of dynamic features, are provided in Appendix A2.1.
Effect of the number of in-context demonstrations. We explore how performance varies with the number of in-context demonstrations using the Qwen2.5-72B-Instruct model. Figure 3 shows that performance improves rapidly as the number increases from 1 to 12, after which it plateaus. This suggests that a moderate number of relevant demonstrations is critical. Too many may introduce irrelevant noise and diminish gains.
Comparison of LLM backbones. Table 4 compares the performance of X-ICM (7B) when replacing Qwen2.5-7B-Instruct with alternative 7B/8B models, including Deepseek-R1-Distill-Qwen-7B [91], Llama3.0-8B-Instruct [92], Ministral-8B-Instruct-2410 [93], and InternLM3-8B-Instruct [94]. The results show significant variability in performance, underscoring the importance of choosing a powerful and well-aligned LLM backbone.
5.3 Real-world Experiments
To evaluate real-world applicability, we test X-ICM on five physical manipulation tasks using an xArm7 robot arm, DH-Robotics gripper, and a third-person Orbbec camera: put block into bin, push button, put bottle into box, stack blocks, and stack cups. We collect 20 demonstrations for each of the five tasks. During testing, we use the demonstrations from the other four tasks to construct the cross-task in-context prompt, enabling zero-shot generalization. The Qwen2.5-72B-Instruct LLM is used as the backbone and the number of seen demonstrations is set to $18$. Each task is executed 20 times, and we report average success rates. Results shown at the bottom of Figure 4 demonstrate the strong real-world cross-task zero-shot performance of our approach. We provide additional details in Section A3 of the Appendix, and showcase some testing videos that include both successful and failed cases in the Supplementary Materials.
Figure 4: Results of five real-world tasks. The tests are conducted in a cross-task zero-shot manner.
6 Conclusions and Discussions
Conclusions. In this work, we have introduced AGNOSTOS, a benchmark for evaluating cross-task zero-shot generalization in robotic manipulation. With 23 unseen tasks across two difficulty levels, AGNOSTOS provides a rigorous testbed for assessing the limits of Vision-Language-Action (VLA) models. Our evaluation of diverse VLA models reveals their significant limitations in unseen task generalization. To address this, we propose X-ICM, a cross-task in-context manipulation method that leverages the cross-task generalization capabilities of LLMs. The X-ICM achieves substantial improvements over existing VLA models, demonstrating robust cross-task zero-shot generalization.
Discussions. While X-ICM significantly enhances cross-task zero-shot generalization, its performance on many unseen tasks, particularly those with both novel objects and motion primitives, remains limited due to LLMs’ challenges in extrapolating beyond pre-training data and in-context demonstrations. In addition, the use of visual information is limited to textualizing object information, which may ignore important visual context in the raw data. To address these limitations, future work could integrate multi-modal reasoning, combining vision, language, and action data to improve generalization to novel semantics. Additionally, leveraging generalizable concepts like object trajectories as a bridge between MLLMs’ perception and dynamic action sequence prediction could reduce cross-task generalization complexity. Scaling X-ICM to diverse robotic embodiments would further enhance its versatility across varied platforms and sensor configurations. We believe the proposed X-ICM benchmark and X-ICM method will inspire future research in generalizable robotic manipulation, paving the way for robots that can seamlessly adapt to open-world environments.

### v3

@@
-Exploring the Limits of Vision-Language-Action Manipulations in Cross-task Generalization
+Exploring the Limits of Vision-Language-Action Manipulation in Cross-task Generalization
 Abstract
 […]
-Figure 1: The proposed AGNOSTOS benchmark evaluates cross-task zero-shot generalization through two difficulty levels.
+Figure 1: The proposed AGNOSTOS benchmark evaluates zero-shot cross-task generalization through two difficulty levels.
 Level-1 testing involves 13 unseen tasks sharing partial similarity (objects or motions) with seen tasks.
@@
 Through large-scale pre-training on diverse data, including human videos [17, 18, 19], real or simulated cross-embodiment robotic demonstrations [3, 20, 21], VLA models can effectively generalize across visual variations within known tasks (i.e., within-task generalization of seen tasks), such as handling objects in novel scenes or with altered properties.
-However, the true promise of VLAs lies in their capacity to generalize across tasks: to handle previously unseen combinations of objects, goals, and actions without prior exposure.
-This capability-cross-task zero-shot generalization-is essential for real-world deployment, where robots are expected to tackle novel tasks as they arise dynamically.
+However, the true promise of VLA models lies in their capacity to generalize across tasks: to handle previously unseen combinations of objects, goals, and actions without prior exposure.
+This capability-zero-shot cross-task generalization-is essential for real-world deployment, where robots are expected to tackle novel tasks as they arise dynamically.
 Despite the rapid progress in VLA research, most prior work [3, 4, 7, 8, 1] has focused on generalization testing in real-world environments.
@@
 Recently, many works [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] have been devoted to developing comprehensive simulated benchmarks, which could be utilized to evaluate the generalization of VLA models.
-While promising, these efforts mainly assess within-task generalization, leaving cross-task zero-shot generalization largely unexplored.
-To address this critical gap, we present AGNOSTOS, a novel benchmark for evaluating cross-task zero-shot generalization in robotic manipulation.
+While promising, these efforts mainly assess within-task generalization, leaving zero-shot cross-task generalization largely unexplored.
+To address this critical gap, we present AGNOSTOS, a novel benchmark for evaluating zero-shot cross-task generalization in robotic manipulation.
 Built on RLBench [25], our benchmark comprises 23 unseen tasks that are carefully curated to differ from 18 commonly used seen training tasks [34, 35].
@@
 Our empirical findings reveal a key limitation: none of the existing models generalize effectively to unseen tasks, highlighting the need for approaches that directly address cross-task generalization.
-Motivated by the success of in-context learning [42, 43, 44] in large language models (LLMs), we propose a Cross-task In-context Manipulation (X-ICM) method to address the challenges of cross-task zero-shot generalization.
+Motivated by the success of in-context learning [42, 43, 44] in large language models (LLMs), we propose a Cross-task In-context Manipulation (X-ICM) method to address the challenges of zero-shot cross-task generalization.
 X-ICM uses demonstrations from seen tasks as in-context examples to prompt LLMs to generate action plans for unseen tasks.
@@
 Our contributions are threefold:
-  • We introduce AGNOSTOS, the first benchmark to systematically evaluate cross-task zero-shot generalization in robotic manipulation for VLAs, with 23 unseen tasks spanning two levels of generalization difficulty.
-  • We propose X-ICM, a novel method that combines in-context prompting and dynamics-guided sample selection to enhance cross-task zero-shot transfer in VLA models.
+  • We introduce AGNOSTOS, the first benchmark to systematically evaluate zero-shot cross-task generalization in robotic manipulation for VLA models, with 23 unseen tasks spanning two levels of generalization difficulty.
+  • We propose X-ICM, a novel method that combines in-context prompting and dynamics-guided sample selection to enhance zero-shot cross-task transfer in VLA models.
   • We conduct extensive evaluations of diverse VLA models on AGNOSTOS, uncovering fundamental limitations in current approaches and demonstrating the superior generalization of X-ICM.
@@
 Evaluating the generalization capabilities of VLA models requires robust and comprehensive benchmarks.
-While existing VLAs mainly focus on real-world testing, which suffers from reproducibility issues and rarely focuses systematically on zero-shot generalization to unseen tasks.
+While existing VLA models mainly focus on real-world testing, which suffers from reproducibility issues and rarely focuses systematically on zero-shot generalization to unseen tasks.
 Recently, many simulated benchmarks [22, 23, 24, 25, 26, 27, 28, 30, 31, 32, 33] have been proposed to offer a controlled environment for rigorous evaluation.
@@
 We find that these benchmarks have predominantly concentrated on evaluating visual variations within tasks, assessing how well models can generalize to new scenes or altered attributes of objects within the same task.
-However, these benchmarks do not focus on the more challenging aspect of cross-task zero-shot evaluation, where models must generalize to entirely new tasks with unseen combinations of object categories and motions.
+However, these benchmarks do not focus on the more challenging aspect of zero-shot cross-task evaluation, where models must generalize to entirely new tasks with unseen combinations of object categories and motions.
 This gap in evaluation limits our understanding of the true generalization capabilities of VLA models.
-Motivated by this, this work proposes the first cross-task zero-shot manipulation benchmark, expanding the evaluation scope of the generalization capabilities of VLAs.
+Motivated by this, this work proposes the first zero-shot cross-task manipulation benchmark, expanding the evaluation scope of the generalization capabilities of VLA models.
 In-context Learning with Large Language Models (LLMs).
@@
 Prior works like KAT [72] and RoboPrompt [76] have shown that off-the-shelf LLMs can predict robot actions directly using within-task in-context samples.
-By extending their paradigms to a cross-task setting, this work specifically focuses on the cross-task zero-shot generalization problem.
+By extending their paradigms to a cross-task setting, this work specifically focuses on the zero-shot cross-task generalization problem.
 Our X-ICM model uses demonstrations from seen tasks as in-context examples to prompt LLMs for action generation in completely unseen tasks.
-Under the cross-task zero-shot setting, the selection of in-context samples [78, 79, 80, 81] is crucial for a robust generalization.
+Under the zero-shot cross-task setting, the selection of in-context samples [78, 79, 80, 81] is crucial for robust generalization.
 Our X-ICM method addresses this challenge by introducing a dynamics-guided sample selection strategy, ensuring that the selected seen demonstrations are relevant to the unseen tasks, thereby stimulating the cross-task generalization capabilities of LLMs.
-3 AGNOSTOS: Cross-task Zero-shot Generalization Benchmark
-To systematically assess the cross-task zero-shot generalization ability of vision-language-action (VLA) models, we introduce AGNOSTOS, a reproducible benchmark built upon the RLBench simulation environment.
+3 AGNOSTOS: Zero-shot Cross-task Generalization Benchmark
+To systematically assess the zero-shot cross-task generalization ability of vision-language-action (VLA) models, we introduce AGNOSTOS, a reproducible benchmark built upon the RLBench simulation environment.
 AGNOSTOS features 18 seen tasks for training and 23 unseen tasks for rigorous generalization testing.
@@
 The comparison evaluates whether each benchmark includes cross-task testing, the types of VLA models evaluated, and the number of Level-1 and Level-2 unseen tasks included.
-Benchmark Simulator No. of train tasks No. of test tasks Cross-task Zero-shot Generalization Test
+Benchmark Simulator
+No. of
+train
+tasks
+No. of
+test
+tasks
+Cross-task Zero-shot Generalization Test
 Cross task Evaluated VLA models Level-1 unseen tasks Level-2 unseen tasks
@@
 Table 1 compares AGNOSTOS with existing robotic manipulation benchmarks in terms of their support for cross-task evaluation, i.e., considering test tasks that have different object categories or motions from training tasks.
-While many benchmarks focus on within-task visual generalization, few explicitly support cross-task zero-shot generalization test, and even fewer include tasks as challenging as our Level-2 testing scenarios.
+While many benchmarks focus on within-task visual generalization, few explicitly support zero-shot cross-task generalization test, and even fewer include tasks as challenging as our Level-2 testing scenarios.
 Even benchmarks that include some form of cross-task evaluation [^2] We carefully categorize their test tasks into our defined two-level difficulty task sets.
-often test on a narrow set of tasks, and typically evaluate only in-domain models, neglecting foundation and human-video pre-trained VLAs.
+often test on a narrow set of tasks, and typically evaluate only in-domain models, neglecting foundation and human-video pre-trained VLA models.
 AGNOSTOS fills this gap by supporting broad model coverage and introducing novel, difficult tasks in Level-2, providing a more comprehensive and diagnostic assessment of cross-task generalization.
 4 X-ICM: Cross-task In-context Manipulation Method
-To push the boundaries of cross-task zero-shot generalization in vision-language-action (VLA) models, we propose a method called Cross-task In-context Manipulation (X-ICM).
+To push the boundaries of zero-shot cross-task generalization in vision-language-action (VLA) models, we propose a method called Cross-task In-context Manipulation (X-ICM).
 Leveraging the cross-task generalization capabilities of LLMs, X-ICM utilizes demonstrations from seen tasks as in-context examples.
 The dynamic characteristics of these examples are used to prompt the LLM to predict action sequences for unseen tasks.
-In contrast to prior works [76, 72] that apply LLMs’ in-context learning to within-task generalization, our work is the first to extend this paradigm to a cross-task zero-shot setting.
+In contrast to prior works [76, 72] that apply LLMs’ in-context learning to within-task generalization, our work is the first to extend this paradigm to a zero-shot cross-task setting.
 A central challenge in this setting is that the selection of in-context demonstrations significantly affects generalization performance.
@@
 4.1 Problem Definition and Method Overview
-We tackle the problem of cross-task zero-shot robotic manipulation by exploiting the in-context learning ability of off-the-shelf LLMs.
-We assume access to a dataset of demonstrations from seen tasks, denoted as $\mathcal{D}^{s}=\{V_{i}^{s},A_{i}^{s},L_{i}^{s}\}_{i=1}^{N}$, where $N$ is the total number of seen demonstrations, $V^{s}_{i}=\{v^{s}_{i,1},\cdots,v^{s}_{i,T}\}$, $A^{s}_{i}=\{a^{s}_{i,1},\cdots,a^{s}_{i,T}\}$, and $L^{s}_{i}$ are the visual observation sequence, action sequence, and language description of the $i$-th seen demonstration ($T$ is the sequence length, which varies for each demonstration).
-For an unseen task with the given initial visual observation $v^{u}$, and language description $L^{u}$, our cross-task zero-shot manipulation method can be formulated as follows:
-$A^{u}_{pred}=\{a^{u}_{1},\cdots,a_{t}^{u}\}=LLM(\mathcal{P}(\mathcal{F}^{sel}(\mathcal{D}^{s}),v^{u},L^{u})),$ (1)
+We tackle the problem of zero-shot cross-task robotic manipulation by exploiting the in-context learning ability of off-the-shelf LLMs.
+We assume access to a dataset of demonstrations from seen tasks, denoted as $\mathcal{D}^{s}=\{V_{i}^{s},A_{i}^{s},L_{i}^{s}\}_{i=1}^{N}$, where $N$ is the total number of seen demonstrations, $V^{s}_{i}=\{v^{s}_{i,1},\@cdots,v^{s}_{i,T}\}$, $A^{s}_{i}=\{a^{s}_{i,1},\@cdots,a^{s}_{i,T}\}$, and $L^{s}_{i}$ are the visual observation sequence, action sequence, and language description of the $i$-th seen demonstration ($T$ is the sequence length, which varies for each demonstration).
+For an unseen task with the given initial visual observation $v^{u}$, and language description $L^{u}$, our zero-shot cross-task manipulation method can be formulated as follows:
+$A^{u}_{pred}=\{a^{u}_{1},\@cdots,a_{t}^{u}\}=LLM(\mathcal{P}(\mathcal{F}^{sel}(\mathcal{D}^{s}),v^{u},L^{u})),$ (1)
 where $\mathcal{F}^{sel}$ is the cross-task in-context demonstration selection process, $\mathcal{P}$ is the cross-task in-context prompt construction process, and $A^{u}_{pred}$ is the predicted action sequence for the tested unseen task.
@@
 The model $\mathcal{G}$ is initialized from InstructPix2Pix [88] and is optimized by:
-$\min_{\mathcal{G}}\mathbb{E}_{i,z,\epsilon\sim\mathcal{N}(0,I)}\left[\left\|\epsilon-\epsilon_{\mathcal{G}}(v^{s}_{i,T,z},z,v^{s}_{i,1},L^{s}_{i})\right\|_{2}^{2}\right],$ (2)
+[…]
 where $z$ is the diffusion timestep, $\epsilon$ is the noise, $v^{s}_{i,T,z}$ is the noised final observation, and $\epsilon_{\mathcal{G}}$ is the noise predictor in model $\mathcal{G}$.
@@
 Dynamics-guided retrieving stage. Once trained, we use $\mathcal{G}$ to extract a dynamic feature $f^{s}_{i}$ for the $i$-th seen demonstration:
-$f^{s}_{i}=[f^{s,vis}_{i},f^{s,lang}_{i}]=\mathcal{G}(v^{s}_{i,1},L^{s}_{i})\in\mathbb{R}^{2\times 1024},$ (3)
+[…]
 where $f^{s,vis}_{i}\in\mathbb{R}^{1024}$ is the predicted latent feature of the target visual observation, $f^{s,lang}_{i}\in\mathbb{R}^{1024}$ is the textual feature of the language description.
@@
 Then, we compute cosine similarities between features and select the top-$K$ seen demonstrations with the highest similarity:
-$R_{i}$ $=\frac{f^{u}\cdot f^{s}_{i}}{\|f^{u}\|\|f^{s}_{i}\|}\quad i\in\{1,\dots,N\},$ (4)
+[…]
 $\mathcal{I}_{idx}$ $=\operatorname{arg\,topk}_{i\in\{1,\dots,N\}}^{K}(R_{i}),$ (5)
@@
 For a tested unseen task, we use the obtained $K$ most-relevant seen demonstrations to construct the in-context prompt for cross-task action prediction with LLMs.
-Our Cross-task In-context Prediction module follows the practices in existing within-task in-context manipulation methods [76, 72], but we are the first to extend this paradigm to the cross-task zero-shot setting.
+Our Cross-task In-context Prediction module follows the practices in existing within-task in-context manipulation methods [76, 72], but we are the first to extend this paradigm to the zero-shot cross-task setting.
 Textualizing key information.
@@
 Additionally, we evenly divide the workspace into $100\times 100\times 100$ grids, and then normalize the 3D coordinates of the objects’ centers into the grid’s coordinate system.
-Thus, for the $m$-th object $O_{k}^{m}$ in the $k$-th seen demonstration, we textualize it as $O_{k}^{m}=$$objectname$: [$x$, $y$, $z$]” (e.g., “block: [52, 50, 19]”).
+Thus, for the $m$-th object $O_{k}^{m}$ in the $k$-th seen demonstration, we textualize it as $O_{k}^{m}=$objectnameobjectname: [$x$, $y$, zz]” (e.g., “block: [52, 50, 19]”).
 For the key-action sequence, following [90], we select the actions when the gripper state changes or the joint velocities are near zero.
@@
 And the angles of end-effector are discretized into 72 bins, with each occupying 5 degrees.
-Thus, for the $t$-th key-action $A_{k}^{t}$ in the $k$-th seen demonstration, we textualize it as $A_{k}^{t}=$ “[$x$, $y$, $z$, $roll$, $pitch$, $yaw$, $gripper$]” (e.g., “[53, 57, 17, 0, 36, 53, 0]”).
+Thus, for the $t$-th key-action $A_{k}^{t}$ in the $k$-th seen demonstration, we textualize it as $A_{k}^{t}=$ “[$x$, $y$, $z$, rollroll, pitchpitch, yawyaw, grippergripper]” (e.g., “[53, 57, 17, 0, 36, 53, 0]”).
 Constructing in-context prompt.
@@
 We textualize all $K$ selected seen demonstrations in this manner, and concatenate them in descending order based on their dynamics similarity scores to the tested unseen task.
-For the tested unseen task, we textualize it using its language description and object context, i.e., “$[L^{unseen},\{O^{1},\cdots,O^{m},\cdots\}]\to$”.
+For the tested unseen task, we textualize it using its language description and object context, i.e., “($L^{unseen},\{O^{1},\@cdots,O^{m},\@cdots\}$$\to$”.
 Finally, the cross-task in-context prompt is formed by concatenating the system prompt, the textualized all seen demonstrations, and the textualized tested unseen task (see the bottom of Figure 2 and an example in Figure A6 of the Appendix).
@@
 5.1 Benchmarking Vision-Language-Action Models
-Table 2 presents the cross-task zero-shot performance of our X-ICM models on 23 unseen tasks from the benchmark, including 13 Level-1 and 10 Level-2 tasks.
+Table 2 presents the zero-shot cross-task performance of our X-ICM models on 23 unseen tasks from the benchmark, including 13 Level-1 and 10 Level-2 tasks.
 These tasks are designed to evaluate the generalization ability of VLA models under varying levels of difficulty.
@@
 Each task is executed 20 times, and we report average success rates.
-Results shown at the bottom of Figure 4 demonstrate the strong real-world cross-task zero-shot performance of our approach.
+Results shown at the bottom of Figure 4 demonstrate the strong real-world zero-shot cross-task performance of our approach.
 We provide additional details in Section A3 of the Appendix, and showcase some testing videos that include both successful and failed cases in the Supplementary Materials.
-Figure 4: Results of five real-world tasks. The tests are conducted in a cross-task zero-shot manner.
+Figure 4: Results of five real-world tasks. The tests are conducted in a zero-shot cross-task manner.
+In addition, we also evaluate the proposed X-ICM model on long-horizon tasks under the zero-shot cross-task setting.
+We use tasks put block into bin, push button, put bottle into box, stack blocks, and pull out the middle block as seen tasks, where we have 10 demonstrations for each task.
+The long-horizon task we evaluated is clean the table, according the robot’s visual observation, we decompose the long-horizon task into three sequential sub-tasks, ie, stack cups, put the stacked cups into plate, place the mango on top of the stacked cups.
+We evaluated this task over 20 rollouts, with sub-task success contingent on the success of all prior sub-tasks.
+For our X-ICM method, the success rates on sub-tasks "stack cups", "put the stacked cups into plate", "place the mango on top of the stacked cups." are 40%, 25%, and 5%, respectively.
+Thus, the overall success rate for the long-horizon task is 5%.
+These results align with expectations, as the sequential nature of long-horizon tasks leads to cumulative failure.
+The zero-shot cross-task setting exacerbates these difficulties.
+These findings highlight the need for more advanced algorithms to handle long-horizon tasks in zero-shot scenarios.
 6 Conclusions and Discussions
 Conclusions.
-In this work, we have introduced AGNOSTOS, a benchmark for evaluating cross-task zero-shot generalization in robotic manipulation.
+In this work, we have introduced AGNOSTOS, a benchmark for evaluating zero-shot cross-task generalization in robotic manipulation.
 With 23 unseen tasks across two difficulty levels, AGNOSTOS provides a rigorous testbed for assessing the limits of Vision-Language-Action (VLA) models.
@@
 To address this, we propose X-ICM, a cross-task in-context manipulation method that leverages the cross-task generalization capabilities of LLMs.
-The X-ICM achieves substantial improvements over existing VLA models, demonstrating robust cross-task zero-shot generalization.
+The X-ICM achieves substantial improvements over existing VLA models, demonstrating robust zero-shot cross-task generalization.
 Discussions.
-While X-ICM significantly enhances cross-task zero-shot generalization, its performance on many unseen tasks, particularly those with both novel objects and motion primitives, remains limited due to LLMs’ challenges in extrapolating beyond pre-training data and in-context demonstrations.
+While X-ICM significantly enhances zero-shot cross-task generalization, its performance on many unseen tasks, particularly those with both novel objects and motion primitives, remains limited due to LLMs’ challenges in extrapolating beyond pre-training data and in-context demonstrations.
 In addition, the use of visual information is limited to textualizing object information, which may ignore important visual context in the raw data.
@@
 We believe the proposed X-ICM benchmark and X-ICM method will inspire future research in generalizable robotic manipulation, paving the way for robots that can seamlessly adapt to open-world environments.
+Acknowledgements.
+This work was supported by the National Natural Science Foundation of China (No.
+62306257), the Guangzhou Municipal Science and Technology Project (No.
+2024A03J0619 and No.
+2024A04J4390) and the Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2023A03J0008), Education Bureau of Guangzhou Municipality.